# GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification

### v3

@@ 6 Results
We evaluated GMLM on three widely recognized heterophilic WebKB benchmark datasets [31]: Cornell, Wisconsin, and Texas. These datasets, involving the classification of university web pages (e.g., student, faculty), are particularly suitable for assessing models on heterophilic graphs (see Appendix A, Table 3 for details).
Our evaluation protocol involved 10 random 60%/20%/20% train/validation/test splits per dataset, reporting mean test accuracy and standard deviation for reproducibility. Baseline models were configured with their reported optimal hyperparameters. GMLM was evaluated with two PLM backbones (GTE-Small and DistilBERT), following the two-stage training procedure from Section 5. Table 1 compares the performance of GMLM variants against a suite of GNN baselines, highlighting the top score in bold and the second highest with an underline.
6.1 Performance on Cornell
On the Cornell dataset, GMLM demonstrates exceptional performance. Specifically, GMLM(DistilBERT) achieves the highest mean test accuracy of 96.87% $\pm$ 0.57, while GMLM(GTE-Small) secures the second-highest accuracy at 96.74% $\pm$ 0.72. Both variants significantly outperform all other baselines. The next best performing baseline is BernNet, which attains 92.13% $\pm$ 1.64, followed by GPRGNN at 91.36% $\pm$ 0.70. Compared to BernNet, GMLM(DistilBERT) shows an improvement of over 4.7%, highlighting GMLM’s superior ability to integrate textual and structural information on this heterophilic graph. The low standard deviations for both GMLM variants indicate stable performance across different random splits.
6.2 Performance on Wisconsin
On the Wisconsin dataset, the highest accuracy is achieved by MLP-2 at 93.87% $\pm$ 3.33, with GPRGNN closely following at 93.75% $\pm$ 2.37. Our GMLM(DistilBERT) variant scores 92.32% $\pm$ 2.21, ranking as the third-best performing model overall, while GMLM(GTE-Small) achieves 91.54% $\pm$ 2.23. Although GMLM does not claim the top position on this dataset, both variants outperform a majority of the other heterophilic GNN baselines, such as APPNP (92.00% $\pm$ 3.59) and H2GCN (87.5% $\pm$ 1.77). The performance of GMLM(DistilBERT) is competitive, being only 1.55% behind the top-performing MLP-2, showcasing its robustness even when simpler models perform strongly.
6.3 Performance on Texas
On the Texas dataset, GMLM again demonstrates state-of-the-art performance. GMLM(DistilBERT) achieves the highest mean test accuracy of 95.16% $\pm$ 1.27, and GMLM(GTE-Small) follows as the second-highest with 94.83% $\pm$ 1.31. Both GMLM variants substantially outperform all other baselines. The next best baseline is BernNet, which attains 93.12% $\pm$ 0.65, followed by GPRGNN at 92.92% $\pm$ 0.61. GMLM(DistilBERT) surpasses BernNet by over 2%. These results underscore GMLM’s effectiveness in handling heterophilic graph structures on the Texas dataset, with consistent performance indicated by their low standard deviations.
7 Timing Comparisons
Table 2 highlights the computational efficiency of our GMLM framework compared to a naive strategy that encodes all node texts at every training iteration with GTE-Small. Across all datasets, GMLM achieves consistent training time reductions of approximately 33% to 40%, attributable to its dynamic active node selection strategy. Additionally, GMLM exhibits marginally lower peak GPU memory usage, with savings of around 6% to 7%, validating its scalability advantage without compromising classification performance.
Dataset Method Pretrain (s) Fine-tune (s) Peak Mem (GB)
Cornell GMLM 1.436 3.081 1.25
Naïve 2.738 3.997 1.33
Texas GMLM 1.534 3.252 1.26
Naïve 2.883 4.012 1.34
Wisconsin GMLM 1.612 3.447 1.26
Naïve 2.994 4.102 1.35
Table 2: Comparison of GMLM vs. a naïve PLM+GNN integration on runtime (s/epoch) and GPU memory usage (GB).
@@ 8 Conclusion
This paper presents GMLM (Graph Masked Language Model), an innovative architecture aimed at seamlessly and effectively merging graph structural details with comprehensive textual meanings for node representation learning, especially in heterophilic graphs. GMLM tackles major obstacles in this area through three primary contributions: (i) a dynamic active node selection strategy that lessens the computational load of handling large text datasets by selectively activating the PLM; (ii) a GNN-specific contrastive pretraining phase employing soft masking with a trainable graph [MASK] token, which prepares the GNN for strong structural feature extraction; and (iii) a specialized fusion network that adaptively merges the generated GNN and PLM embeddings. Our comprehensive tests on popular heterophilic benchmark datasets (Cornell, Texas, and Wisconsin) confirm the effectiveness of GMLM. The findings indicate that GMLM, especially when employing PLMs such as DistilBERT and GTE-Small, attains either state-of-the-art or significantly competitive performance against a wide range of established GNN baselines. For example, GMLM(DistilBERT) achieved peak accuracies of 96.87% on Cornell and 95.16% on Texas, greatly surpassing previous approaches. Even with the Wisconsin dataset, where a basic MLP variant excelled, GMLM stayed competitive, highlighting its resilience. The reliable outcomes over several attempts and the small standard deviations emphasize the robustness of our method.
9 Limitations
The GMLM architecture, by design, utilizes textual information associated with graph nodes. This introduces a dependency on the nature of the input dataset, which constitutes the primary limitation of the current study. The effectiveness of GMLM is contingent upon the availability of raw textual data for the nodes. This work targets graphs with node-level raw textual content, which is essential for leveraging the PLM component in GMLM. Our choice of the WebKB datasets (Cornell, Texas, Wisconsin) is deliberate, in many widely used citation or product graphs (e.g., Cora, Citeseer, PubMed), node features are limited to bag-of-words vectors or preprocessed embeddings, and the original raw text is unavailable. Without node text, GMLM reduces to a pure GNN.
@@ Appendix A Dataset Statistics
Table 3 provides an overview of the key statistics for the datasets used in our experiments.
Name Nodes Edges Features Classes
Cornell 183 298 1,703 5
Texas 183 325 1,703 5
Wisconsin 251 515 1,703 5
@@
Appendix B Model Sizes
Table 4 details the approximate number of trainable parameters for the main components of our GMLM architecture. The RGCN branch size is calculated based on the configuration used for the WebKB datasets (Cornell, Texas, Wisconsin) which have an input feature dimension of 1,703 and a GNN hidden channel size of 512.
Component Parameters (Approx.)
GTE-Small (PLM) 33M
DistilBERT (PLM) 66M
RGCN Branch (GNN) $\sim$18.88M
Calculated for GNN input features $d_{x}=1703$, hidden channels $d_{h}=512$, 3 RGCN layers,
   and ‘num_relations=4‘. Includes GNN layers, batch norms, residual projections, and
   the GNN-specific mask token.
Table 4: Approximate model sizes (number of trainable parameters).
Appendix C Hyperparameter Ablation
This section provides further details on the hyperparameter tuning process for our GMLM architecture and an ablation analysis of key components. To identify the optimal configuration, we performed a grid search over several critical hyperparameters. The specific parameter ranges and values explored during this search are detailed in Table 5.
Figure 3 visualizes the performance (mean test-set accuracy across 10 runs using DistilBERT as the PLM backbone) of various hyperparameter configurations using a parallel coordinates plot. Each line in the plot represents a unique combination of hyperparameters, allowing for an intuitive understanding of how different settings influence the final accuracy on the benchmark datasets (Cornell, Wisconsin, and Texas). The axes correspond to the hyperparameters: soft masking interpolation weight ($\beta$), the range for active node selection ratio ($p_{active\_nodes}$), the range for GNN input perturbation ratio ($p_{gnn\_mask}$), and the contrastive pretraining temperature ($\tau_{pretrain}$). The plot highlights the configuration that yielded the best overall performance, as indicated by the thicker orange line and specified in the legend as (0.7, 0.55, 0.35, 0.7), where 0.55 and 0.35 represent the midpoints of the selected ranges for $p_{active\_nodes}$ (0.3, 0.8) and $p_{gnn\_mask}$ (0.2, 0.5) respectively. This visualization aids in understanding the sensitivity of GMLM to these hyperparameters and validates the choices made for the final reported results.
Appendix D Impact of Training Stages
To qualitatively assess the impact of different training stages on the learned node representations, Figure 2 visualizes the evolution of fused node embeddings for the Cornell dataset, using DistilBERT as the PLM backbone. The figure presents t-SNE (left column) and PCA (right column) projections of these embeddings at three key stages: before any training, after the GNN contrastive pretraining stage, and after complete end-to-end fine-tuning. Node colors in the plots correspond to their respective ground-truth classes.
Figure 2: Evolution of node embeddings on the Cornell dataset (DistilBERT PLM) visualized using t-SNE (left column) and PCA (right column) at different training stages: (Top Row) Before Training, (Middle Row) After GNN Contrastive Pretraining, and (Bottom Row) After Complete End-to-End Fine-tuning. Colors represent distinct node classes (0-4). The variance captured by the first two principal components is noted for PCA plots (e.g., "Var: 0.34").
Before Training (Top Row): Initially, before any training, the t-SNE plot shows that node embeddings from different classes are largely intermingled and lack clear separation. Similarly, the PCA visualization (Var: 0.34) indicates that the initial features, likely dominated by the raw input features and randomly initialized GNN/PLM components not yet fine-tuned, capture limited class-specific variance in the principal components. There is no discernible clustering according to node labels.
After Pretraining (Middle Row): Following the GNN contrastive pretraining stage, we observe some emergent structuring in the embedding space. Both t-SNE and PCA plots show early signs of clustering, with some same-colored points beginning to group together. This suggests that the contrastive pretraining has started to instill structurally relevant information into the GNN embeddings, leading to initial differentiation. However, the class separation is still not well-defined, and considerable overlap remains. The PCA plot (Var: 0.40) shows a slight increase in the explained variance by the top two components, aligning with this initial structural learning.
After Complete Training (Bottom Row): After the end-to-end fine-tuning phase, where both the GNN and the DistilBERT PLM are trained jointly for the node classification task, there is a marked improvement in class separability. The t-SNE visualization clearly shows more distinct and compact clusters for the different node classes. The PCA plot (Var: 0.61) now captures a substantially higher proportion of the variance in its first two principal components, reflecting the enhanced class structure in the embedding space. This indicates that the GMLM model, through its two-stage training and fusion mechanism, has learned to produce highly discriminative embeddings that are well-suited for the downstream node classification task.
Figure 3: Mean test-set accuracy across 10 runs for different hyperparameter configurations using DisitilBERT as the PLM. The averaged values such as 0.55 and 0.35 refer to (0.3,0.8) and (0.2,0.5) respectively.
Soft Mask $\boldsymbol{\beta}$ Active Node Ratio GNN Perturb. Ratio Contrastive Temp. $\boldsymbol{\tau}$
0.5 (0.2, 0.6) (0.1, 0.3) 0.3
0.7 (0.3, 0.8) (0.2, 0.5) 0.5
0.9 (0.4, 0.9) (0.3, 0.6) 0.7
Table 5: Grid search parameters explored for GMLM hyperparameter tuning. Ranges for ratio-based parameters are specified as (min, max) for uniform sampling during training.
Appendix E Method Comparison with other integrations of PLMs and GNNs
To further contextualize our proposed GMLM architecture, this section provides a detailed comparative analysis against prominent existing methods in graph-based neural network pre-training and integration with language models. Table 6 (from the main paper, re-iterated here for completeness or referenced if preferred) summarizes the key characteristics, objectives, and methodological differences between GMLM and influential works such as GPT-GNN [18], GraSAME [13], and the Graph Language Model (GLM) framework by Plenz and Frank [12]. This comparison highlights the unique aspects of GMLM, including its specific GNN contrastive pre-training strategy, the dynamic active node selection for efficient PLM engagement, and the explicit fusion mechanism for combining structural and textual representations, particularly for the task of heterophilic node classification.
Feature GPT-GNN GraSAME GLM GMLM (Ours)
Primary Goal Generative pre-training of GNNs for transfer learning. Inject token-level graph structure into PLMs for graph-to-text generation. Enable LMs to natively process graphs & text as Graph Transformers. Integrate GNN (structure) & PLM (rich node text) for heterophilic node classification.
Pre-training Focus GNN pre-training (generative: attribute & edge reconstruction). Relies on pre-trained PLM. GNN component trained during fine-tuning (PLM frozen). Initializes from pre-trained LM; explores LM’s adaptability to graph structure. GNN contrastive pre-training (soft masking, learnable [MASK] token). PLM is pre-trained & then fine-tuned.
GNN Component Any standard GNN (e.g., HGT, GCN). GNN (SAGE, GAT, RGCN) operates on token-level graph derived from linearized input. LM architecture itself is modified to become a Graph Transformer (GT). Dedicated RGCN branch operating on the input graph structure.
PLM Component Not a primary focus for rich node text; can use simpler models (e.g., LSTM) for text attributes. Standard PLM (e.g., T5 encoder), parameters are frozen during fine-tuning. LM (e.g., T5) architecture is modified to become a GT; parameters initialized from PLM. Standard PLM (e.g., GTE-Small, DistilBERT) for node text; parameters are fine-tuned.
Integration/ Fusion Method GNN learns attribute-structure dependency implicitly via the generative pre-training task. GNN output directly guides the Q vector in PLM’s self-attention mechanism. Transformer self-attention & positional encoding are adapted for graph processing within the LM architecture. GNN & PLM embeddings are explicitly concatenated & fused via a dedicated MLP.
Text/Attribute Handling Focuses on "node attributes" (can be vectors or generated text). Graph is linearized to text; GNN processes a token-level graph derived from this. Triplet components (textual) are processed by the GT-adapted LM. PLM processes raw textual descriptions of selected nodes. GNN uses original node features (potentially perturbed).
Scalability Strategy Subgraph sampling and an adaptive queue for large-scale graph pre-training. Reduces trainable parameters significantly by freezing the PLM. Not a primary focus of the presented work. Dynamic active node selection for PLM processing and loss computation.
Table 6: Comparison of GMLM with related language model integrations with graph neural networks. GMLM distinguishes itself through its specific GNN contrastive pre-training, dynamic active node selection for PLM engagement, and explicit fusion of separately processed GNN and PLM embeddings for node classification.
Appendix F LLM Usage

### v4

@@ 6 Results
We evaluate GMLM on five widely-used heterophilic benchmark datasets. Following standard protocols, we randomly split the nodes into train, validation, and test sets with a proportion of 48%/32%/20% for the Cornell, Texas, Wisconsin, and Actor datasets, and 50%/25%/25% for Amazon following Platonov et al. (2024). All reported results in Table 1 are the mean accuracy (%) and standard deviation averaged over 10 independent runs.
As detailed in the table, our GMLM framework achieves state-of-the-art performance on the majority of benchmarks. The GMLM-Snowflake-embed variant establishes new records on four of the five datasets: Cornell, Texas, Wisconsin, and Amazon. The performance gains are particularly striking on the Texas dataset, where our model achieves 97.18% accuracy, surpassing the strongest LLM-based baseline by over 8 absolute percentage points. Similarly, it improves the state-of-the-art on Wisconsin by nearly 5 points to 92.04%.
Crucially, GMLM consistently outperforms both specialized heterophily-GNNs and even the much larger Vicuna 7B model. This demonstrates that our architectural design, which deeply integrates a pre-trained GNN with a PLM via bi-directional cross-attention, is more effective and capital-efficient than simply applying a larger, general-purpose model. On the Actor dataset, our result is highly competitive, nearly matching the 7B model’s performance with a significantly smaller PLM backbone.
A comparison of our model variants reveals that performance scales with the representational power of the PLM, as Snowflake-embed (305M) consistently outperforms GTE-Base (110M). However, even the smaller GMLM-GTE-Base is highly competitive, outperforming the majority of all baselines and securing the second-best result on Wisconsin. This confirms the robustness and effectiveness of the core GMLM framework itself.
@@ 7 Conclusion
In this paper, we introduced the Graph Masked Language Model (GMLM), a novel framework that effectively and efficiently integrates deep graph structural learning with the rich semantic understanding of Pre-trained Language Models. By employing a two-stage training procedure featuring a contrastive GNN pre-training phase and an end-to-end fine-tuning stage with dynamic active node selection, our model addresses key challenges of scalability and robust representation learning. The use of a sophisticated bi-directional cross-attention module for fusion proved critical, enabling GMLM to achieve new state-of-the-art results on four out of five widely-used heterophilic benchmarks. Our model significantly outperforms both specialized GNNs and much larger LLM-based approaches, underscoring that a carefully designed, deeply integrated architecture can be more powerful and capital-efficient than brute-force scaling for text-rich graph representation learning.
8 Limitations and Future Work
Despite the strong performance of GMLM, we acknowledge the limitations that present opportunities for future research. The framework’s effectiveness is intrinsically tied to the availability and quality of textual descriptions for nodes. In scenarios where such texts are sparse, noisy, or entirely absent, the PLM branch offers diminished value. Future work could explore robust fallbacks or methods for generating informative pseudo-texts to enhance performance on text-poor graphs.
@@ Appendix A Dataset Statistics
Table 2 provides an overview of the key statistics for the datasets used in our experiments from HeTGB Li et al. (2025).
Name Nodes Edges H($G$) Classes
Cornell 195 304 0.13 5
Texas 187 328 012 5
Wisconsin 265 530 0.20 5
Actor 4416 12172 0.56 5
Wisconsin 24492 93050 0.38 5
@@
Appendix B LLM Usage